### v1

@@ 3.5 Basin Computation with Validated Gradient Flow
return $D(M)$
Ascending basins for local minima follow the same structure with reversed inequality conditions: we accept neighbors $v$ with $\hat{y}(v)>\hat{y}(u)$ and require $\text{cm}(\hat{y};P(u,v))\geq\theta$ for long edges. The complete partition of vertices into basins emerges from running this algorithm for all detected local extrema, with ambiguous vertices (reachable from multiple extrema) assigned through tie-breaking rules based on path length and function value differences.
3.6 Computational Complexity and Practical Implementation
The computational cost of validated basin computation scales with the number of long edges requiring validation. Let $n=|V|$ denote the number of vertices, $k$ the neighborhood size in the k-nearest neighbor graph, and $L$ the number of long edges exceeding threshold $\tau_{q}$. For each long edge, computing the shortest alternative path requires breadth-first search with cost $O(nk)$ in the worst case, though typical paths have length $O(k)$ leading to expected cost $O(k^{2})$ per validation. With $L$ long edges, total validation cost is $O(Lk^{2})$.
Since $L$ is typically a small fraction of total edges (for $q=0.75$, we have $L=0.25\cdot nk$), and most validations occur during basin construction where we only examine edges incident to basin vertices, the practical cost is substantially lower than the worst-case bound. We implement several optimizations to improve efficiency further.
First, we cache shortest paths between frequently accessed vertex pairs. When multiple extrema have overlapping neighborhoods, the same alternative paths arise repeatedly during basin computation. A least-recently-used cache with capacity $O(n)$ reduces redundant path computations at the cost of $O(n)$ additional memory.
Second, we relax the strict monotonicity requirement through the threshold $\theta$, avoiding validation failure due to minor numerical fluctuations in the smoothed function $\hat{y}$. Setting $\theta=0.90$ accommodates up to 10% of the total variation in non-monotonic differences while maintaining robust filtering of artifactual edges.
Third, we defer validation to the breadth-first search phase rather than pre-computing validity for all edges. This lazy evaluation ensures we validate only edges that actually participate in basin computation, reducing unnecessary work for edges far from extrema.
The edge length quantile $\tau_{q}$ provides a tunable parameter balancing computational cost and geometric accuracy. Smaller $q$ validates fewer edges (faster computation, more aggressive filtering), while larger $q$ validates more edges (slower computation, more conservative filtering). Empirically, $q\in[0.75,0.85]$ provides good performance across diverse datasets, with the optimal value depending on the intrinsic dimensionality of the data manifold and the smoothness of the outcome function.
3.7 Return to the Motivating Example
We return to the two-crescent example to demonstrate how path monotonicity validation resolves basin jumping. Recall that naive gradient flow produces trajectories jumping between crescents due to long edges connecting their tips. Under validated gradient flow, we examine each such long edge $e=[i,j]$ where $i$ belongs to the first crescent and $j$ to the second.
Computing the shortest alternative path $P(i,j)$ yields a route that must travel along one crescent toward the connecting region, cross to the other crescent, then proceed to the target. For a function with separate minima in each crescent, this path necessarily exhibits non-monotonic behavior: the function decreases from $i$ toward the minimum in the first crescent, increases through the connecting region, then decreases again toward $j$ and its local minimum. The path co-monotonicity coefficient $\text{cm}(\hat{y};P(i,j))$ captures this oscillation, yielding a value near zero rather than the $\theta$ required for validation.
The validation procedure rejects these artifactual long edges, forcing gradient trajectories to remain within their respective crescents. Each local minimum now correctly captures all vertices in its crescent, and the basin boundaries align with the geometric separation between regions. The resulting Morse-Smale complex accurately reflects the intrinsic structure of the outcome landscape rather than artifacts of the discrete graph construction.
This path-based approach provides direct enforcement of the defining property of gradient trajectories while remaining computationally practical. The validated basins serve as geometric stratification for subsequent inference, enabling regional analysis that respects the manifold structure underlying the data. We proceed to develop co-monotonicity measures that quantify feature associations within these geometrically faithful regions.
@@ 4.1 The Inadequacy of Global Association Measures
Mutual information and related entropy-based measures address nonlinearity but require density estimation in high dimensions, introducing computational expense and methodological complexity. For large graphs with thousands of vertices, estimating joint distributions becomes impractical. We require computationally efficient measures that capture directional relationships while respecting graph geometry.
@@ 4.2 Edge-Based Directional Concordance
To illustrate these concepts, consider a one-dimensional example where vertices lie on a line and edges connect adjacent points. Let $y(v)$ represent temperature at position $v$, and $z(v)$ represent humidity. If temperature and humidity both increase as we move along the line (both functions have positive derivatives), then $\Delta_{e}y\cdot\Delta_{e}z>0$ for all edges $e$, reflecting strong positive co-monotonicity. If temperature increases while humidity decreases, the products are negative, indicating negative co-monotonicity. If temperature increases in some regions and decreases in others, while humidity varies independently, the products will have mixed signs, and we require aggregation over neighborhoods to assess overall association.
@@ 4.3 Vertex-Level Co-Monotonicity Coefficients
This vertex-level formulation provides spatially-resolved information about where in the graph the two functions are most strongly associated. Unlike global correlation, which reduces all vertex pairs to a single number, co-monotonicity coefficients reveal regional structure in association patterns. In microbiome applications, this enables identification of samples where particular phylotypes exhibit strong co-variation with outcomes, while other samples show weak or opposite associations.
@@ 4.4 Weighting Schemes and Geometric Interpretations
4.5 Connection to Angular Correlation of Gradients
@@
We provide an informal argument connecting derivative-weighted co-monotonicity to geometric quantities on smooth manifolds. Suppose data $X=\{x_{1},\ldots,x_{n}\}$ are sampled from a smooth Riemannian manifold $M$, and let $y,z:M\to\mathbb{R}$ be smooth functions. The k-nearest neighbor graph $G_{n}$ constructed from $X$ approximates $M$ in the limit of dense sampling with appropriate choice of $k=k(n)$.
@@
For a vertex $v$ corresponding to point $x_{v}\in M$, the neighborhood $N(v)$ consists of points approximately uniformly distributed in a geodesic ball of radius $r_{k}$ around $x_{v}$, where $r_{k}$ is the distance to the $k$-th nearest neighbor. The edge difference $\Delta_{e}y/\Delta_{e}$ for an edge $e=[v,u]$ approximates the directional derivative of $y$ in the direction from $x_{v}$ to $x_{u}$:
$\frac{\Delta_{e}y}{\Delta_{e}}\approx\langle\nabla_{M}y(x_{v}),\xi_{e}\rangle,$ (25)
where $\xi_{e}$ is the unit tangent vector at $x_{v}$ pointing toward $x_{u}$, and $\nabla_{M}y$ is the Riemannian gradient of $y$ on $M$. Similarly, $\Delta_{e}z/\Delta_{e}\approx\langle\nabla_{M}z(x_{v}),\xi_{e}\rangle$.
The numerator of $\text{cm}_{\partial}(y,z)(v)$ becomes a Riemann sum approximating an integral over directions:
$\sum_{u\in N(v)}\frac{\Delta_{e}y}{\Delta_{e}}\cdot\frac{\Delta_{e}z}{\Delta_{e}}\approx\int_{S_{x_{v}}}\langle\nabla_{M}y(x_{v}),\xi\rangle\langle\nabla_{M}z(x_{v}),\xi\rangle\,d\sigma(\xi),$ (26)
@@
where $S_{x_{v}}$ is the unit sphere in the tangent space $T_{x_{v}}M$ and $d\sigma$ is the uniform measure. Evaluating this integral yields
$\int_{S_{x_{v}}}\langle\nabla_{M}y(x_{v}),\xi\rangle\langle\nabla_{M}z(x_{v}),\xi\rangle\,d\sigma(\xi)=C_{d}\langle\nabla_{M}y(x_{v}),\nabla_{M}z(x_{v})\rangle,$ (27)
where $C_{d}>0$ is a dimension-dependent constant arising from integrating products of directional projections over the sphere. The denominator undergoes similar approximation, leading to
@@
$\lim_{n\to\infty}\text{cm}_{\partial}(y,z)(v)=\frac{\langle\nabla_{M}y(x_{v}),\nabla_{M}z(x_{v})\rangle}{|\nabla_{M}y(x_{v})|\cdot|\nabla_{M}z(x_{v})|},$ (28)
the cosine of the angle between the gradients of $y$ and $z$ at point $x_{v}$.
@@
This limiting interpretation reveals that derivative-weighted co-monotonicity measures the directional alignment of gradients in the underlying continuous geometry. Functions with parallel gradients (angle near zero) have co-monotonicity near +1+1, functions with anti-parallel gradients (angle near $\pi$) have co-monotonicity near 1-1, and functions with orthogonal gradients (angle near $\pi/2$) have co-monotonicity near zero. This geometric perspective explains why derivative weighting is natural for smooth functions on manifolds: it captures intrinsic geometric association independent of the coordinate system used to represent the data.

### v2

@@ 3.5 Basin Computation with Validated Gradient Flow
return $D(M)$, dist, pred
Ascending basins for local minima follow the same algorithmic structure with reversed monotonicity conditions. To construct $A(m)=\{v:\pi_{\downarrow}(v)=m\}$ for a local minimum $m$, we initialize with $A(m)=\{m\}$ and perform breadth-first exploration accepting neighbors $v$ of vertices $u\in A(m)$ when $\hat{y}(v)>\hat{y}(u)$. This explores upward from $m$ (in the ascending direction) while tracing descending flow trajectories backward, identifying all vertices from which descending flow reaches $m$. For long edges, we require $\text{cm}(\hat{y};\gamma_{[u,v]})\geq\theta$, validating that the alternative path from $u$ to $v$ exhibits ascending monotonicity consistent with the upward exploration direction.
The complete partition of vertices into basins emerges from running this algorithm for all detected local extrema. Due to the non-uniqueness discussed above, some vertices may be discovered during exploration from multiple extrema of the same type. We resolve such conflicts through tie-breaking rules based on discovery order: the first extremum to reach a vertex during breadth-first exploration claims that vertex for its basin. This deterministic rule ensures each vertex belongs to exactly one ascending basin and exactly one descending basin, though the specific assignment depends on the order in which extrema are processed. Alternative tie-breaking strategies include assigning ambiguous vertices based on minimal path length to extrema, maximal function value difference along paths, or probabilistic assignment proportional to path validation scores. The sensitivity of boundary vertex assignments to these algorithmic choices reinforces that robust inference should focus on the interior of basins rather than their boundaries.
For vertices that remain unassigned after processing all extrema, we have two options. The first is to leave them as isolated singletons, which may occur for vertices in flat regions where no clear gradient direction exists. The second is to perform a final assignment phase where unassigned vertices are allocated to the nearest basin based on graph distance, function value similarity, or other proximity measures. In practice, validated gradient flow with appropriate quantile and monotonicity thresholds produces nearly complete basin coverage, with unassigned vertices arising primarily in highly ambiguous flat regions or poorly connected boundary zones.
Computational Complexity and Practical Considerations
The algorithm’s computational cost decomposes into several components. The initial quantile computation requires sorting all edge lengths, an $O(|E|\log|E|)$ operation performed once before basin construction begins. For each extremum, the breadth-first exploration visits each vertex at most once and examines each incident edge at most once, giving $O(|V|+|E|)$ per extremum in the absence of validation. When validation is required for a long edge $e=[u,v]$, computing the shortest alternative path using breadth-first search costs $O(|E|)$ in sparse graphs. Evaluating the co-monotonicity coefficient along a path of length $\ell$ costs $O(\ell)$, typically $O(\log|V|)$ for paths in graphs with good expansion properties.
The total validation cost depends on the quantile threshold $q$ and the graph geometry. Choosing $q\in[0.75,0.90]$ ensures that only the longest 1010-2525% of edges require validation. In well-structured graphs where long edges are rare and localized, validation overhead remains modest. In pathological cases where many long edges connect distant regions, validation costs can approach $O(L\cdot|E|)$ where $L$ is the number of long edges examined during all basin explorations. For typical applications with n=500n=500-$5000$ vertices and k=10k=10-$50$ nearest neighbors, basin computation completes in seconds to minutes on modern hardware, with validation adding at most a factor of 22-55 overhead compared to naive gradient flow.
Memory requirements are dominated by storing the graph structure ($O(|E|)$ for adjacency lists and edge lengths), the distance and predecessor maps for each basin ($O(|V|)$ per extremum), and temporary storage for shortest path computations ($O(|V|)$). For graphs with $n=10^{4}$ vertices and moderate degree, total memory consumption remains well under 1 GB, enabling in-memory computation on standard workstations. Parallel computation across extrema is straightforward, as basin constructions are independent until the final tie-breaking phase, offering near-linear speedup on multi-core systems.
@@ 4.1 The Inadequacy of Global Association Measures
Mutual information and related entropy-based measures address nonlinearity and can be computed on graph structures, providing viable alternatives for quantifying associations between vertex functions. However, these measures produce a single global summary of the relationship between $y$ and $z$ across all vertices. They cannot reveal spatial heterogeneity in association patterns, nor can they identify regions where the relationship is strong versus weak, or positive versus negative. When association structure varies across the graph, as commonly occurs in gradient flow cells with different monotonic behaviors, a single global measure obscures this regional variation.
Co-monotonicity coefficients address this limitation by providing vertex-level measures rather than global summaries. For each vertex $v$, we compute a coefficient $c(y,z;v)$ that quantifies directional concordance between $y$ and $z$ within the local neighborhood of $v$. This yields a function $c:V\to[-1,1]$ that maps vertices to association strengths, enabling identification of regions where features exhibit strong positive co-variation, regions with negative relationships, and regions where associations are weak or absent. The coefficients have intuitive geometric interpretation: they measure whether $y$ and $z$ tend to increase or decrease together across edges incident to each vertex, directly capturing the directional concordance that characterizes monotonic co-variation within gradient flow structures. This vertex-level resolution proves essential for regional inference, where statistical power comes from identifying coherent association patterns within cells while allowing relationships to differ across cells.
@@ 4.4 Weighting Schemes and Geometric Interpretations
4.5 Alternative Normalization and Geometric Interpretation
The co-monotonicity coefficient defined in Section 4.3 uses absolute-value normalization to measure the proportion of concordant versus discordant directional changes across edges. We now introduce a complementary measure with Pearson-style normalization that weights edges by the magnitudes of their changes. This variant admits a rigorous geometric interpretation as the cosine of the angle between gradient vectors on smooth manifolds, providing theoretical justification for the derivative-weighted form. The two normalizations capture related but distinct aspects of directional concordance, and their differences prove informative for understanding local association structure.
Correlation-Type Co-Monotonicity Coefficient
We begin with the alternative definition. Given functions $y,z:V\to\mathbb{R}$ on a weighted graph $G=(V,E)$ with edge weights $w_{e}\geq 0$, the correlation-type co-monotonicity coefficient at vertex $v$ is
$\text{cm}_{\text{cor}}(y,z;w)(v)=\frac{\sum_{u\in N(v)}w_{e}\Delta_{e}y\cdot\Delta_{e}z}{\sqrt{\sum_{u\in N(v)}w_{e}(\Delta_{e}y)^{2}}\sqrt{\sum_{u\in N(v)}w_{e}(\Delta_{e}z)^{2}}},$ (16)
where $e=[v,u]$ denotes the edge connecting $v$ and $u$, and $\Delta_{e}y=y(u)-y(v)$, $\Delta_{e}z=z(u)-z(v)$ are the edge differences. When either denominator term vanishes (indicating that one of the functions is constant across all edges incident to $v$), we define $\text{cm}_{\text{cor}}(y,z;w)(v)=0$ by convention.
The normalization by geometric mean of squared changes ensures that $\text{cm}_{\text{cor}}(y,z;w)(v)\in[-1,1]$ for all vertices, with extremal values achieved under perfect proportionality. If $\Delta_{e}z=k\Delta_{e}y$ for some constant $k>0$ across all edges incident to $v$, then $\text{cm}_{\text{cor}}(y,z;w)(v)=1$. If $\Delta_{e}z=-k\Delta_{e}y$ for $k>0$, then $\text{cm}_{\text{cor}}(y,z;w)(v)=-1$. The measure vanishes when positive and negative products balance in a squared-magnitude-weighted sense, which differs subtly from the absolute-value-weighted balance in the definition of Section 4.3.
For derivative-weighted co-monotonicity where $w_{e}=1/\Delta_{e}$ with $\Delta_{e}$ denoting edge length, the coefficient takes the form
$\text{cm}_{\text{cor},\partial}(y,z)(v)=\frac{\sum_{u\in N(v)}\frac{\Delta_{e}y}{\Delta_{e}}\cdot\frac{\Delta_{e}z}{\Delta_{e}}}{\sqrt{\sum_{u\in N(v)}\left(\frac{\Delta_{e}y}{\Delta_{e}}\right)^{2}}\sqrt{\sum_{u\in N(v)}\left(\frac{\Delta_{e}z}{\Delta_{e}}\right)^{2}}},$ (17)
which directly correlates the discrete directional derivatives across the neighborhood of $v$. This form proves essential for the manifold interpretation we develop below.
Geometric Interpretation via Gradient Alignment
We now establish a rigorous connection between correlation-type co-monotonicity and angular correlation of gradients on smooth Riemannian manifolds. This provides both theoretical justification for the derivative-weighted formulation and intuitive geometric meaning for the coefficient values.
@@ 4.5 Alternative Normalization and Geometric Interpretation
Suppose data $X=\{x_{1},\ldots,x_{n}\}$ are sampled from a smooth $d$-dimensional Riemannian manifold $M$ embedded in $\mathbb{R}^{D}$ for $D\geq d$. Let $y,z:M\to\mathbb{R}$ be smooth functions with non-vanishing gradients in a neighborhood of a point $x_{v}\in M$. The k-nearest neighbor graph $G_{n}$ constructed from $X$ approximates $M$ in the limit of dense sampling, with appropriate scaling $k=k(n)$ such that $k/n\to 0$ and $k/\log n\to\infty$ as $n\to\infty$.
@@
For a vertex $v$ corresponding to point $x_{v}\in M$, the neighborhood $N(v)$ consists of the $k$ nearest sample points, which are approximately uniformly distributed in a geodesic ball $B_{r_{k}}(x_{v})$ of radius $r_{k}$ around $x_{v}$, where $r_{k}$ is the distance to the $k$-th nearest neighbor. As $n\to\infty$, we have $r_{k}\to 0$ with rate determined by the local density of the sampling distribution and the scaling of $k$.
Consider an edge $e=[v,u]$ connecting $v$ to a neighbor $u$ at point $x_{u}\in M$ with $d_{M}(x_{v},x_{u})=\Delta_{e}$, where $d_{M}$ denotes the Riemannian distance on $M$. For small $\Delta_{e}$, the function difference can be approximated by the directional derivative:
$\Delta_{e}y=y(x_{u})-y(x_{v})=\langle\nabla_{M}y(x_{v}),\xi_{e}\rangle\Delta_{e}+O(\Delta_{e}^{2}),$ (18)
where $\xi_{e}\in T_{x_{v}}M$ is the unit tangent vector at $x_{v}$ pointing in the direction of the geodesic from $x_{v}$ to $x_{u}$, and $\nabla_{M}y$ denotes the Riemannian gradient of $y$ on $M$. Dividing by edge length gives
$\frac{\Delta_{e}y}{\Delta_{e}}=\langle\nabla_{M}y(x_{v}),\xi_{e}\rangle+O(\Delta_{e}).$ (19)
Similarly, for the second function,
$\frac{\Delta_{e}z}{\Delta_{e}}=\langle\nabla_{M}z(x_{v}),\xi_{e}\rangle+O(\Delta_{e}).$ (20)
The numerator of $\text{cm}_{\text{cor},\partial}(y,z)(v)$ aggregates products of these discrete directional derivatives over the neighborhood $N(v)$. As $n\to\infty$ and the neighborhood becomes dense, this sum approximates a Riemann integral over directions. Specifically, for uniformly distributed neighbors in a geodesic ball, the directions $\{\xi_{e}:u\in N(v)\}$ become equidistributed on the unit sphere $S^{d-1}\subset T_{x_{v}}M$ in the tangent space. We obtain
$\sum_{u\in N(v)}\frac{\Delta_{e}y}{\Delta_{e}}\cdot\frac{\Delta_{e}z}{\Delta_{e}}$ $\approx\sum_{u\in N(v)}\langle\nabla_{M}y(x_{v}),\xi_{e}\rangle\langle\nabla_{M}z(x_{v}),\xi_{e}\rangle$ (21)
$\to|N(v)|\int_{S^{d-1}}\langle\nabla_{M}y(x_{v}),\xi\rangle\langle\nabla_{M}z(x_{v}),\xi\rangle\,d\sigma(\xi),$ (22)
@@
where $d\sigma$ denotes the uniform probability measure on the unit sphere $S^{d-1}$.
We evaluate this integral using a standard result from directional statistics. For any two vectors $\mathbf{a},\mathbf{b}\in\mathbb{R}^{d}$, the integral of their directional projections squared over the unit sphere satisfies
$\int_{S^{d-1}}\langle\mathbf{a},\xi\rangle\langle\mathbf{b},\xi\rangle\,d\sigma(\xi)=\frac{1}{d}\langle\mathbf{a},\mathbf{b}\rangle.$ (23)
This identity follows from expanding the inner products in coordinates and using the fact that for $i\neq j$, we have $\int_{S^{d-1}}\xi_{i}\xi_{j}\,d\sigma(\xi)=0$ by symmetry, while $\int_{S^{d-1}}\xi_{i}^{2}\,d\sigma(\xi)=1/d$ by normalization (since $\sum_{i=1}^{d}\int_{S^{d-1}}\xi_{i}^{2}\,d\sigma(\xi)=\int_{S^{d-1}}|\xi|^{2}\,d\sigma(\xi)=1$ and all $d$ components contribute equally). Applying this to $\mathbf{a}=\nabla_{M}y(x_{v})$ and $\mathbf{b}=\nabla_{M}z(x_{v})$, we obtain
$\sum_{u\in N(v)}\frac{\Delta_{e}y}{\Delta_{e}}\cdot\frac{\Delta_{e}z}{\Delta_{e}}\to\frac{|N(v)|}{d}\langle\nabla_{M}y(x_{v}),\nabla_{M}z(x_{v})\rangle.$ (24)
The denominator terms undergo similar approximation. We have
$\sum_{u\in N(v)}\left(\frac{\Delta_{e}y}{\Delta_{e}}\right)^{2}$ $\approx\sum_{u\in N(v)}\langle\nabla_{M}y(x_{v}),\xi_{e}\rangle^{2}$ (25)
$\to|N(v)|\int_{S^{d-1}}\langle\nabla_{M}y(x_{v}),\xi\rangle^{2}\,d\sigma(\xi)$ (26)
$=\frac{|N(v)|}{d}|\nabla_{M}y(x_{v})|^{2},$ (27)
where the final equality again uses the directional integral formula with $\mathbf{a}=\mathbf{b}=\nabla_{M}y(x_{v})$. Similarly,
$\sum_{u\in N(v)}\left(\frac{\Delta_{e}z}{\Delta_{e}}\right)^{2}\to\frac{|N(v)|}{d}|\nabla_{M}z(x_{v})|^{2}.$ (28)
Combining these results, we obtain the limiting formula:
@@
$\lim_{n\to\infty}\text{cm}_{\text{cor},\partial}(y,z)(v)$ $=\frac{\frac{|N(v)|}{d}\langle\nabla_{M}y(x_{v}),\nabla_{M}z(x_{v})\rangle}{\sqrt{\frac{|N(v)|}{d}|\nabla_{M}y(x_{v})|^{2}}\sqrt{\frac{|N(v)|}{d}|\nabla_{M}z(x_{v})|^{2}}}$ (29)
$=\frac{\langle\nabla_{M}y(x_{v}),\nabla_{M}z(x_{v})\rangle}{|\nabla_{M}y(x_{v})|\cdot|\nabla_{M}z(x_{v})|}$ (30)
$=\cos\theta(x_{v}),$ (31)
where $\theta(x_{v})\in[0,\pi]$ denotes the angle between the gradient vectors $\nabla_{M}y(x_{v})$ and $\nabla_{M}z(x_{v})$ in the tangent space $T_{x_{v}}M$.
@@
This limiting interpretation provides compelling geometric meaning for the correlation-type co-monotonicity coefficient. When $\text{cm}_{\text{cor},\partial}(y,z)(v)\approx 1$, the gradients of $y$ and $z$ are nearly parallel at the corresponding point in $M$, indicating that both functions increase in approximately the same direction. When $\text{cm}_{\text{cor},\partial}(y,z)(v)\approx-1$, the gradients are anti-parallel, so the functions increase in opposite directions. When $\text{cm}_{\text{cor},\partial}(y,z)(v)\approx 0$, the gradients are approximately orthogonal, indicating that the functions vary independently in different tangent directions. This geometric perspective explains why derivative weighting is natural for smooth functions on manifolds: it captures intrinsic geometric association that is independent of the ambient coordinate system used to represent the data.
Comparison of Normalization Strategies
We now examine how the two normalization strategies differ and when each is preferable for statistical applications. Denote the absolute-value coefficient from Section 4.3 as $\text{cm}_{\text{abs}}(y,z;w)(v)$ and the correlation-type coefficient as $\text{cm}_{\text{cor}}(y,z;w)(v)$. Both have numerators $\sum_{u\in N(v)}w_{e}\Delta_{e}y\cdot\Delta_{e}z$, but different denominators.
The measures answer related but distinct questions about directional concordance. The absolute-value coefficient asks: ”What proportion of directional changes are concordant versus discordant?” It computes a signed sum of weighted products and normalizes by the total absolute magnitude of those products. Each edge contributes to the denominator according to $|w_{e}\Delta_{e}y\cdot\Delta_{e}z|$, which depends on both the weight and the product magnitude. Edges where both functions change substantially receive more weight in the denominator, but the absolute value prevents cancellation between positive and negative products. The resulting measure can be interpreted as the net directional concordance relative to the total amount of joint variation, treating edges democratically in the sense that the normalization accounts for how much each edge contributes to the concordance signal.
The correlation-type coefficient asks: ”How strongly are the directional derivatives correlated?” It normalizes by the geometric mean of squared changes, so the denominator is $\sqrt{\sum w_{e}(\Delta_{e}y)^{2}}\sqrt{\sum w_{e}(\Delta_{e}z)^{2}}$. This treats the functions separately before combining: we measure the total variation of each function independently, then correlate their directional changes. Edges where one or both functions vary substantially contribute more to their respective denominator terms, making the measure sensitive to the magnitude profile of changes. This aligns with classical Pearson correlation, which measures linear association weighted by the variances of the two variables.
By the Cauchy-Schwarz inequality applied to the sequences $(a_{u}=\sqrt{w_{e}}\Delta_{e}y)_{u\in N(v)}$ and $(b_{u}=\sqrt{w_{e}}\Delta_{e}z)_{u\in N(v)}$, we have
$\sum_{u\in N(v)}w_{e}|\Delta_{e}y\cdot\Delta_{e}z|\leq\sqrt{\sum_{u\in N(v)}w_{e}(\Delta_{e}y)^{2}}\sqrt{\sum_{u\in N(v)}w_{e}(\Delta_{e}z)^{2}},$ (32)
with equality if and only if $|\Delta_{e}y|$ and $|\Delta_{e}z|$ are proportional across all edges: $|\Delta_{e}z|=k|\Delta_{e}y|$ for some constant $k\geq 0$ independent of $u$. When the numerator is positive (indicating net concordance), dividing both sides by the numerator and taking reciprocals shows that
$|\text{cm}_{\text{cor}}(y,z;w)(v)|\leq|\text{cm}_{\text{abs}}(y,z;w)(v)|,$ (33)
with equality when changes are proportional in magnitude across edges. The absolute-value coefficient generally has larger magnitude because its denominator is smaller (or equal), reflecting that it requires less total variation to achieve a given level of net concordance.
The measures coincide in specific situations that reveal when normalization choice is inconsequential. Perfect proportionality $\Delta_{e}z=k\Delta_{e}y$ for constant $k$ across all edges incident to $v$ produces $\text{cm}_{\text{abs}}=\text{cm}_{\text{cor}}=\text{sign}(k)$, since both denominators equal $|k|$ times the same quantity. When all edge weights are equal and all magnitudes $|\Delta_{e}y|$ and $|\Delta_{e}z|$ are approximately constant across edges, the two normalizations differ only by a constant factor that cancels in the ratio. These scenarios represent homogeneous neighborhoods where all edges provide comparable information about directional concordance.
The measures diverge most strongly when magnitude heterogeneity is substantial. Consider a vertex $v$ with two neighbors $u_{1}$ and $u_{2}$, where $\Delta_{e_{1}}y=0.1$, $\Delta_{e_{1}}z=0.1$, while $\Delta_{e_{2}}y=0.1$, $\Delta_{e_{2}}z=5$, with unit edge weights. Both edges show concordance (positive products), but the second edge exhibits much larger magnitude change in $z$. The absolute-value coefficient computes
$\text{cm}_{\text{abs}}=\frac{0.01+0.5}{0.01+0.5}=1,$ (34)
indicating perfect concordance since both edges agree in direction. The correlation-type coefficient gives
$\text{cm}_{\text{cor}}=\frac{0.01+0.5}{\sqrt{0.01+0.01}\sqrt{0.01+25}}=\frac{0.51}{\sqrt{0.02}\sqrt{25.01}}\approx 0.228,$ (35)
a much smaller value reflecting that the large change in $z$ on the second edge is not matched by a correspondingly large change in $y$, indicating imperfect correlation of magnitudes despite consistent direction. The absolute-value measure emphasizes directional agreement, while the correlation measure penalizes magnitude mismatch.
For statistical applications, the choice between normalizations depends on the inferential goals and data characteristics. We recommend using $\text{cm}_{\text{abs}}$ when robustness to occasional extreme edges is desired. In noisy data where individual edges may exhibit large but unreliable changes due to measurement error or sampling artifacts, the absolute-value normalization prevents such edges from dominating the association measure. The frequency-based interpretation (proportion of concordant changes) remains meaningful even when individual change magnitudes are uncertain. This normalization is also preferable when the scientific question focuses on consistency of directional relationships rather than strength of linear association.
The correlation-type $\text{cm}_{\text{cor}}$ is preferable when edges represent measurements of varying precision and larger changes carry more information. In applications where the magnitude of change reflects confidence or biological significance, weighting by squared magnitudes appropriately up-weights reliable signals. The geometric interpretation as cosine of gradient angles provides theoretical grounding and connects to classical differential geometry. When downstream analyses assume correlation-like structure (such as principal component analysis or factor models applied to vertex-level coefficients), using $\text{cm}_{\text{cor}}$ ensures compatibility with those methods.
In practice, computing both measures and examining their differences provides diagnostic information about local association structure. Vertices where $|\text{cm}_{\text{abs}}|\gg|\text{cm}_{\text{cor}}|$ indicate regions where directional concordance is consistent but magnitude matching is poor, suggesting that the association may be monotonic but non-linear. Vertices where the two measures are comparable suggest homogeneous neighborhoods with proportional changes, indicating robust local association. Vertices where both measures are small indicate weak or inconsistent relationships regardless of normalization. This dual perspective enriches the analysis by revealing not just whether functions co-vary but how the pattern of co-variation manifests across neighborhood edges.
The availability of two complementary measures also enables sensitivity analysis for downstream inference. If regional clustering or hypothesis testing produces similar results using either $\text{cm}_{\text{abs}}$ or $\text{cm}_{\text{cor}}$, this provides evidence that the findings are robust to normalization choice and likely reflect genuine biological or scientific patterns rather than artifacts of measurement scale. Conversely, if results differ substantially, this flags vertices or regions where conclusions depend critically on whether we prioritize directional consistency versus magnitude correlation, prompting closer examination of the underlying data and scientific interpretation.